How do you write an equation in slope-intercept form when given the x and y-intercepts?
Example:
x-intercept 3, y-intercept 2

Answer by venugopalramana(3286)

How do you write an equation in slope-intercept form when given the x and y-intercepts?
Example:
x-intercept 3,
MEANS VALUE OF X=3 WHEN Y=0...THAT IS (3,0)IS A POINT ON THE LINE.
y-intercept 2
MEANS VALUE OF Y=2 WHEN X=0....THAT IS (0,2) IS A POINT ON THE LINE.
STD.EQN.IS
Y=MX+C...WHERE M=SLOPE AND C= INTERCEPT
SO IF Y INTERCEPT =2 THEN
2=M*0+C.......C=2
IF X INTERCEPT IS 3 THEN
0=M*3+C=3M+2
3M=-2
M=-2/3
HENCE EQN. OF LINE IS
Y=-2X/3+2